# Statistical Principles for Platform Trials

### v2

@@ Abstract
While within a clinical study there may be multiple doses and endpoints, across different studies each study will result in either an approval or a lack of approval of the drug compound studied. The False Approval Rate (FAR) is the proportion of drug compounds that lack efficacy incorrectly approved by regulators. (In the U.S., compounds that have efficacy and are approved are not involved in the FAR consideration, according to our reading of the relevant U.S. Congressional statute).
@@ 1 An motivating example: the HEALEY ALS Platform trial
Platform trials represent an emerging clinical trial design gaining popularity for their ability to conduct simultaneous testing of different treatments. More significantly, they enable the continuous incorporation of new treatments during the study, facilitating continuous learning. While platform trial design enhances resource efficiency and improves the likelihood of study participants receiving active treatments rather than placebos, the novelty of this design approach leaves ample room for exploration within the field of clinical trials, particularly regarding statistical principles due to the variety of potential outcomes.
While the excitement surrounding platform trials is undeniable, they do come with several limitations. For instance, platform trials are not universally applicable to all diseases, and meticulous planning is necessary to ensure their success. Nevertheless, despite these challenges, there have been instances of successful implementations of platform trials. One such case is the HEALEY ALS Platform trial, sponsored by Massachusetts General Hospital. Amyotrophic lateral sclerosis (ALS), commonly known as Lou Gehrig’s disease, is a progressive and fatal neuromuscular disease with the majority of ALS patients dying within 2 to 5 years of receiving a diagnosis.
@@
In the HEALEY ALS Platform trial, the primary outcome is the ALS Functional Rating Scare-Revised (ALSFRS-R) score, which measures 12 aspects of physical functionality. It consists of 12 questions, each targeting a specific aspect of functional capability, with individual scores ranging from 4 (normal function) to 0 (complete loss of function). Thus, the aggregate score can span from a maximum of 48, indicating full functionality, to a minimum of 0, reflecting total loss of physical function. For a repeatedly measured continuous variable such as ALSFRS-R score, Mixed Models Repeated Measures (MMRM) analyses are typically employed.
@@ 2 False Approval Rate (FAR) in a regulatory setting
For regulatory agencies like the European Medical Agency (EMA) and the United States Food and Drug Administration (FDA), the purpose of Type I error rate control is to control the rate of incorrect regulatory decisions while balancing the benefits and risks to patients. When it is correctly done, Type I error rate control translates to False Approval Rate (FAR) control, i.e., controlling the rate of approving a treatment (e.g., drug) when it is not efficacious, across infinitely many studies (which are not identical, nor necessarily independent). In this paper, we refer to each clinical trail as a “study”.
@@ 2.1 Two layers of families and error rates
Within each study, there may be multiple doses and/or endpoints. The within-study family of inferential statements is the collection of (prespecified) statements to be made about them (such as confidence intervals for low dose efficacy in the primary endpoint and high dose efficacy in the secondary endpoint).
@@
In the Healey ALS platform trial, within each regimen, there exists a family of efficacy assessment on primary outcomes (e.g., ALSFRS-R total score disease progression) and secondary outcomes (e.g., respiratory function). Subsequently the regimens in the platform trial forms a higher layer of family of studies, each of which either succeeds or fails in claiming efficacy of the compound compared to the placebo.
@@ 2.2 Tukey’s three Multiple Comparison error rates
It can perhaps be said that Tukey (1953), which has been reprinted as Tukey (1994), started Multiple Comparisons as a discipline in Statistics. The form of statistical inference that John W. Tukey preferred was confidence intervals. He defined the following error rates:
@@
Therefore, instead of the transposed familywise error rate (abbreviated as FWER), we use Tukey’s original ordering error rate familywise (abbreviating it as ERFw). Also contrary to popular perception, it is not true that the only alternative to ERFw is false discovery rate (FDR). While FDR control may be appropriate for medical image reconstruction, it is well known that FDR control is inappropriate for controlling errors within a clinical trial because it can be manipulated by sponsors (see Section 6 of Finner and Roter 2001). Which multiple comparison error rate to control depends on the specific application as we will discuss below.
2.3 Error Rate per Family (ERpF) control guards regulatory False Approval Rate (FAR)
Within a study (a clinical trial), an incorrect statement about a primary endpoint can lead to incorrect approval of a compound, and an incorrect statement about a secondary endpoint can lead to incorrect claim of additional efficacy on the drug label. Since any incorrect statement can lead to a (regulatory) incorrect decision, this has led to controlling error rate familywise within each study since the early 2000s. We will show that across studies, controlling error rate per family appropriately balances the risk and benefit to patients, regardless of whether the studies are independent or correlated in families of Platform trial studies.
Even though individual studies and those in platform trials are conducted at overlapping times, we can think of them as occurring in a (countably) infinite sequence of clinical trials, indexed as Study $\ldots,-1,0,1,\ldots$. Consider a finite number of $k$ studies, some of which are in platform trials while others are not. Within these $k$ studies, some compounds have efficacy while others do not. Studies with efficacious compounds that fail, failing to reject the null hypothesis of no-efficacy, do not lead to incorrect regulatory approval. We thus isolate those studies with compounds that lack efficacy (and therefore should not gain regulatory approval) as studies with index in $N$, and let $n$ denote its cardinality (the number of studies with index in $N$). To control the error rate per family (ERpF) of studies in $N$ at $5\%\times n$ per a family of $n$ studies means, as $n\rightarrow\infty$, the proportion of compounds lacking efficacy that incorrectly gain regulatory approval approaches $5\%$.
Different treatment arms within a Platform trial may share a common control arm and therefore not independent. Among the infinitely many studies across independent as well as those in Platform trials, the following theorem allows us to show if error rate familywise (ERFw) of every study within each Platform trial is set at the 5% level, this dependence does not affect regulatory false approval rate being no more than 5-per-100 (or 5%), so long as dependence is short term (within each Platform trial only, while studies in different Platform trials remain independent).
@@ 2.3 Error Rate per Family (ERpF) control guards regulatory False Approval Rate (FAR)
Let $\ldots,Y_{-1},Y_{0},Y_{1},\ldots$ be the indicator functions of past, present, and future studies that lead to efficacy claims from rejection of the null hypothesis that the compound studied lacks efficacy. Among studies with index in $N$ that should not lead to regulatory approvals, we relabel their $Y_{i}$ as $X_{i}$. We show that controlling ERFw within each study controls regulatory false approval rate ERpF across studies, regardless of whether these studies are in platform trials or not.
Let $X_{i}$ be the indicator of a Type I error being committed as a result of Study $i,~{}i\in N,$ a study with compound that lacks efficacy,
@@
then $\mu_{i}\leq 0.05$ for all $i\in N$. For $X_{i},X_{j}$ within the same Platform trial, the correlation $\rho_{j-i}$ in (2.1) is between zero and one. But with short term dependence, for $X_{i},X_{j}$ in different Platform trials, $\rho_{j-i}=0$. For example, if the practical maximum number of study arms within a single Platform trial is 100, then the $\rho_{j-i}$ in (2.1) would satisfy
@@
Let $m$ = $|j-i|$, then $\rho_{m}=0,~{}m=101,102,\ldots$ for Platform trials. Therefore the series $\sum_{1}^{\infty}\rho_{i}$ is convergent. The variance convergence condition is also satisfied since $\sigma^{2}_{i}=\mu_{i}(1-\mu_{i})\leq 1$ and $\sum_{i=1}^{\infty}(\log i)^{2}/i^{2}$ is a convergent series. As $\frac{1}{n}\sum_{i=1}^{n}\mu_{i}\leq\frac{1}{n}\sum_{i=1}^{n}0.05=0.05$, the long run regulatory FAR will not exceed 5-per-100 ($5\%$) if the Type I error rate familywise (ERFw) of each study is controlled at $5\%$. Consequently, for a regulatory agency such as the Food and Drug Administration (FDA) or the European Medical Agency (EMA) to not approve medicine that are ineffective more than 5-per-100 across infinitely many studies, controlling the error rate familywise (ERFw) of each study at the 5% level guarantees that, according to Theorem 2.1.
2.4 Additive multiplicity adjustment for regulatory approval
Congress of the United States enacted in 1938 the Food, Drug, & Cosmetic Act. Together with its 1962 Kefauver-Harris Amendment, they require pharmaceutical companies to submit evidence of new drugs’ safety and effectiveness for approval to market to the public. To balance risk-benefit to the public, the U.S. Food and Drug Administration (FDA) has been allowing five incorrect approvals per 100 ineffective compounds, 5-per-100, as follows.
The FDA requires control of the error rate familywise (ERFw) within each study to be at $5\%$ strongly, so that even for a complex study (with multiple doses and endpoints in combination), the chance that an ineffective compound will show efficacy is no more than $5\%$. Then, the regulatory mission is accomplished across any family of $n$ studies with inefficacious compounds because the error rate per family (ERpF) is controlled at 5-per-100 ($5\%\times n$), as follows.
For a family of $n$ studies of inefficacious compounds with index in $N$, let $V$ be the number of studies with Type I errors in this family. Then $V=\sum_{i\in N}X_{i}$ where $X_{i}$ is as in (2.3) the indicator that a Type I error is committed as a result of Study $i,~{}i\in N$. Expectation of a sum of random variables simply add. Suppose $P\{X_{i}=1\}=5\%$ exactly (not conservatively) for all $i,~{}i\in N$. Then
$E[V]$ $=$ $\sum_{i\in N}P(X_{i}=1)=5\%\times n$
exactly, not conservatively, regardless of whether $X_{i},~{}i\in N,$ are independent, positively dependent, or negatively dependent. So in terms of the expected number of incorrect regulatory approvals, controlling the Type I error rate familywise (ERFw) of each study exactly at 5% leads to controlling the incorrect regulatory approval rate at exactly 5% for a family of $n$ studies with inefficacious compounds (which we phrase as 5-per-100). This is as expected of a regulatory agency, balancing risk and benefit to patients, contrary to the perception given by Bai et al (2020).
3 False Discovery Rate (FDR) in a regulatory setting
3.1 FDR would change policy
False Discovery Rate (FDR) control has been proposed for Type I error rate control across studies in Platform trials (add reference here). FDR is the expected proportion of rejections (among rejections both correct and incorrect) that are incorrect [^2] Rejecting a false null hypothesis is “correct”, while rejecting a true null hypothesis is “incorrect”.. Let $H_{0i},i=1,\ldots,k$ be the null hypotheses that each compound in a Platform trial with $k$ studies lacks efficacy, then
$=$ $\mathrm{E}\left[\frac{\mbox{number of inefficacious compounds incorrectly approved }}{\mbox{total number of compounds approved, both efficacious and inefficacious}}\right]$
While current regulatory practice of controlling FAR (which is Tukey’s per family error rate across studies) limits the number of in-efficacious compounds incorrectly approved to 5-per-100, controlling FDR in Platform trials would increase that number to beyond 5-per-100.
Suppose a Platform trial has $n^{\mathrm{true}}$ true no-efficacy nulls $H_{0i}^{{\scriptsize\mbox{true}}},i=1,\ldots,n^{\mathrm{true}}$, the expected number of incorrect rejections (incorrect approvals) is just
$\mathrm{ERpF}$ $=$ $\sum_{v=1}^{n^{\mathrm{true}}}P\left\{\mbox{Reject }H_{0i}^{{\scriptsize\mbox{true}}}\right\}$ (3.2)
Shouldn’t the $v$ here be $i$? So to compute $E[V]$ we add up the Type I error rate of each study. If FDR were fixed, then to do this calculation we can try reverse-engineering $P\left\{\mbox{Reject }H_{0i}^{{\scriptsize\mbox{true}}}\right\}$ and adding them up. For easier subsequent comparison with FDR (which is a fraction), we can divide ERpF by the total number of drug compounds that should not be approved
$\dfrac{\mathrm{EFpF}}{n^{\mathrm{true}}}$ $=$ $\sum_{v=1}^{n^{\mathrm{true}}}\dfrac{v}{n^{\mathrm{true}}}P\{V=v\}$ (3.3)
$\mathrm{FDR}$ $=$ $\sum_{v=1}^{n^{\mathrm{true}}}\dfrac{v}{v+n^{\mathrm{false}}}P\{V=v\}$ (3.4)
$\leq$ $\sum_{v=1}^{n^{\mathrm{true}}}\dfrac{v}{0+n^{\mathrm{false}}}P\{V=v\}.$ (3.5)
How ERpF and FDR count the number of incorrect approvals $v$ differ is evident by comparing (3.4) with (3.3). FDR attenuates the number of incorrect rejection $v$ by the number of correct rejection of false nulls.
From (3.5), we see $n^{\mathrm{false}}$ being not small (relative to $v$) allows inflation of $P\left\{\mbox{Reject }H_{0i}^{{\scriptsize\mbox{true}}}\right\}$ while keeping FDR small. For simplicity, suppose each Platform trial has $k=2$ studies only, and suppose the compound in Study $w$ is highly efficacious but the compound in Study $v$ is in-efficacious:
$FDR$ $=$ $E\left[\frac{V}{R}\right]$
$=$ $\frac{1}{2}\times 10\%=5\%$
Switching to FDR control would thus be a policy change: For every efficacious compound approved, regulators reward the industry with an additional fungible False Approval Rate (FAR) token [^3] We term the tokens fungible because any company can cash them in for any compound. toward approval of inefficacious compounds. We raise this caution as we do not quite see in the Food, Drug, & Cosmetic Act or the Kefauver-Harris Amendment language indicating intent that guarding against incorrectly approving inefficacious drug compounds should be affected by the number of efficacious compounds correctly approved.
3.2 Integrity of error rates against manipulation is desirable
Whereas the False Approval Rate (FAR) is only concerned with the rate at which inefficacious compounds are incorrectly approved, clearly the issue with FDR is its denominator contains both correct and incorrect approvals of compounds. Similar to the concern in Section 6 of Finner and Roter (2001), we wonder whether that might subject FDR to manipulation as follows.
If FDR were to be controlled across studies in Platform trials involving multiple drug companies, seems to us an astute company would not put forth their best compound, leaving that to other companies. Even in Platform trials involving a single sponsor, we wonder whether there would be temptation to include highly efficacious compounds of low commercial value, to benefit compounds with low efficacy but high commercial value.
@@ 4 Stability of FAR and Conditional Assessment
While reassuring that a 5% ERFw within each study results in an ERpF of 5% across all the studies, there is a concern regarding the stability of ERpF across Platform trials.
@@ 4.1 Conditional Stability of FAR is desirable
Section F titled Multiplicity of FDA’s draft guidance on Master Protocols states
An error rates which is an expectation is just the expected value aspect of the probability distribution of errors made. For any unconditional error rate, even with it being controlled at 5%, it could be the case that some families of studies (with particular characteristics of their studies within) can have actual error rates that are rather variable (varying from 1% to 9% say), while other families of studies can have error rates that are more stable (ranging from 4% to 6% only say). So, in addition to the unconditional (expected) error rate, variability of the actual error rate should be considered as well, conditionally given characteristics of studies within families. The specific characteristics of concerns from sharing a control in a platform trial are as follows:
Mean $\hat{\mu}_{SC}$
An under-performing or over-achieving estimated mean $\hat{\mu}_{SC}$ of the shared control makes getting approval currently easier or harder for all the new compounds, with under-performing $\hat{\mu}_{SC}$ especially of concern to regulatory agencies;
Correlation $\hat{\rho}_{SC}$
A high realized (overall) correlation measure $\hat{\rho}_{SC}$ among the test statistics associated with the new compounds can substantially inflate the variability of $V$, the number of incorrect discoveries/approvals. Note that correlation this can be represented as a product of a function of the ratio of sample size $\lambda_{i}=\sqrt{1+\frac{n_{SC}}{n_{i}}}$. Specifically, $\hat{\rho}_{SC}=\lambda_{i}\lambda_{j}$ for all $1<=i<j<k$. These $\lambda$ values can easily be found via existing software. We favor this notation due to the ease of extension to more complicated structures. For the HEALEY ALS pltform tril, $\lambda=$ (0.674, 0.722, 0.671, 0.690) to a close approximation.
To facilitate the assessment of impact of correlations, analogous to commerce on the internet keeping track of cookies, we propose an informational “cookie” for each treatment/regimen of a platform trial to be kept track of.
Control Sharing Matrix
Cookie for each treatment/regimen should include the exact time interval of the enrollment period of each regimen, so that by knowing the “overlap” for the enrollment periods of a pair of regimens, one can compute the pairwise covariance and correlation for the active treatment versus Control comparisons.
$n$ A B C D
A 122 122 164 103
B 122 120 122 103
C 164 122 120 103
D 103 103 103 75
Table 1: Control Sharing Matrix (CSM) for the HEALEY ALS Platform trial (inferred as best we can from Results posted at clinicaltrials.gov): For each regimen, the corresponding diagonal element is the number of patients randomized to the active treatment. For each pair of regimens, the corresponding off-diagonal element is the number of patients that they have in common in the shared Control that each active treatment is compared to.
A B C D
A 1 0.498 0.425 0.478
B 0.498 1 0.496 0.476
C 0.425 0.496 1 0.476
D 0.478 0.476 0.476 1
Table 2: Matrix of correlations between $\hat{\mu}_{i}-\hat{\mu}_{SC_{ij}}$ and $\hat{\mu}_{j}-\hat{\mu}_{SC_{ij}}$
Within each new treatment arm, randomization ratio $n_{Rx}:n_{C}$ of new compound to the shared control can range from a low of $1:1$ (in which case the sample size ratio of the shared control to each new compound becomes $k:1$) to a high of $k:1$ (in which case the sample size ratio of the shared control to each new compound becomes $1:1$). Given a fixed total sample size for the platform trial, choice of the randomization ratio $n_{Rx}:n_{C}$ involves multiple considerations.
Patients would prefer a higher $n_{Rx}:n_{C}$ randomization ratio since higher it is higher the chance of receiving a new treatment compound.
From an individual drug sponsor’s point of view, a benefit of joining a platform trial is pooling control patients allows a randomization ratio higher than the $1:1$ in a standalone study, to more accurately estimate effects of the new compounds. However, that is balanced by the master protocol sponsor’s view that lower the $n_{Rx}:n_{C}$ randomization ratio higher the sample size of the shared control, which lessens the chance that an overachieving shared control sample causes all the new compounds difficulty demonstrating efficacy.
On the other hand, from the regulators’ point of view, stability of ERpF control is affected by both stability of the sample mean of the shared control as well as correlation among the pivotal statistics. A lower randomization ratio increases sample size of the shared control, which lessens the chance of an under-performing shared control sample giving all the new compounds an easy pass. A lower randomization ratio also decreases correlation among the pivotal statistics, which makes ERpF control more stable.
For example, none of the four drugs tested in Regimens ABCD of the HEALEY ALS Platform trial demonstrated efficacy in the primary endpoint ALSFRS-R, [^4] Active treatment for Regimen A was Zilucoplan, B was Verdiperstat, C was CNM-Au8, and D was Pridopidine. which can be due to either the compounds lacking efficacy, or from an overachieving shared control, or both.
Choice of the randomization ratio $n_{Rx}:n_{C}$ should take all these considerations into account. Our app lets one find the sweet spot.
4.2 Effect of mean and correlation induced by sharing control on $V^{\star}$
It is to be noted that while both mean and correlation induced by sharing control on $V^{\star}$ individually affect the distribution of $V$ and thus $V^{\star}$, we need to consider the simultaneous effect on the distribution. The variance of $V$ in particular is of note. As seen in Figure 1, for cases when $\hat{\mu}_{SC}$ is greater than the treatment arm mean $\hat{\mu}_{i}$, the performance of the standard deviation of $V$ is much more stable. On the other hand, when $\hat{\mu}_{SC}$ is less than $\hat{\mu}_{i}$, the behavior of the standard deviation is much more volatile. This tells that in cases like this, it is beneficial to look at the whole behavior of the probability distribution of V rather than single instances.
Figure 1: The behavior of standard deviation for V is heavily affected by correlation and sample control. We assume the true $\mu_{sc}=0$, i.e. it is non efficacious. As we decrease the value of our $\hat{\mu}_{SC}$, the standard deviation of $V$ shows less stability.
$\hat{\mu}_{SC}$ = -3
$\hat{\mu}_{SC}$ = -1.5
$\hat{\mu}_{SC}$ = 0
$\hat{\mu}_{SC}$ = 1.5
$\hat{\mu}_{SC}$ = 3
Figure 2: Snapshot of different standardized $\hat{\mu}_{SC}$ values on correlation’s effect on standard deviation where $\mu_{i}=0$. As can be seen, when the standardized $\hat{\mu}_{SC}$ is less than the treatment mean, the behavior is less simple whereas for $\hat{\mu}_{SC}$ is greater than the treatment mean, the standard deviation of V decreases as correlation increases.
4.3 Effect of correlation induced by sharing control on $V^{\star}$
Let us say there are 20 new treatments in a platform trial, and 10 of the treatments show efficacy. Let $V^{\star}$ denote the number of treatments out of those ten with their confidence intervals for efficacy not covering their true parameter values, then $E[V^{\star}]=0.05\times 10=0.5$. If these new treatment arms were independent, then $StdDev(V^{\star})$, the standard deviation of $V^{\star}$ (under the null) is 0.69 which is larger than the mean $E[V^{\star}]=0.5$. Further, correlation among the test statistics associated with the new compounds (induced by sharing a control) makes the variability of $V^{\star}$ (the number of incorrect approvals) even bigger. In the scenario above, a common correlation of 0.5 among the test statistics increases $StdDev(V^{\star})$ from 0.69 to to 1.16. This ever-present variability of the actual error rate in a particular family of clinical trials is a fact of life that we believe should receive more attention in practice, regardless of the error rate being ERpF or ERFw. We show how this calculation is done below.
@@ 4.3 Effect of correlation induced by sharing control on $V^{\star}$
Let $N$ denote the subset of index $\{1,\ldots,k\}$ of compounds that lack efficacy. For $i\in N$, let $Z_{i}$ be the indicator that the $i^{th}$ confidence interval erroneously show efficacy. Then $V^{\star}=\sum_{i\in N}Z_{i}$. Since $Var(V^{\star})=\sum_{i}Var(Z_{i})+\sum_{i\neq j}Cov(Z_{i},Z_{j})$ and $Var(Z_{i})=\alpha(1-\alpha)$, we just need the covariances $Cov(Z_{i},Z_{j}),i\neq j,$ to compute $Var(V^{\star})$.
There are various ways of computing the non-coverage events covariances among the treatment arms in a platform trial. For proof of concept, consider platform trials in therapeutic areas where it is reasonable to take each pair of $(T^{\star}_{i},T^{\star}_{j}),i\neq j,$ as bivariate standard Normal. Alzheimer’s Disease (AD) and Type 2 Diabetes Mellitus (T2DM) are examples of such therapeutic areas. Then

### v3

@@ Abstract
While within a clinical study there may be multiple doses and endpoints, across different studies each study will result in either an approval or a lack of approval of the drug compound studied. The term False Approval Rate (FAR) is the term this paper utilizes to represent the proportion of drug compounds that lack efficacy incorrectly approved by regulators. (In the U.S., compounds that have efficacy and are approved are not involved in the FAR consideration, according to our reading of the relevant U.S. Congressional statute).
@@ 1 An motivating example: the HEALEY ALS Platform Trial
Platform trials represent an emerging clinical trial design gaining popularity for their ability to conduct simultaneous testing of different treatments. More significantly, they enable the continuous incorporation of new treatments during the study, facilitating continuous learning. While platform trial design enhances resource efficiency and improves the likelihood of study participants receiving active treatments rather than placebos, the novelty of this design approach leaves ample room for exploration within the field of clinical trials, particularly regarding statistical principles due to the variety of potential outcomes (Koenig et al.,, 2023).
While the excitement surrounding platform trials is undeniable, they do come with several limitations. For instance, platform trials are not universally applicable to all diseases. Meticulous planning is also necessary to ensure their success (Koenig et al.,, 2023).. Nevertheless, despite these challenges, there have been instances of successful implementations of platform trials. One such case is the HEALEY ALS Platform Trial, sponsored by Massachusetts General Hospital.
@@
ALS is a progressive and fatal neuromuscular disease with the majority of ALS patients dying within 2 to 5 years of receiving a diagnosis. In the HEALEY ALS Platform Trial, the primary outcome is the ALS Functional Rating Scare-Revised (ALSFRS-R) score, which measures 12 aspects of physical functionality. It consists of 12 questions, each targeting a specific aspect of functional capability, with individual scores ranging from 4 (normal function) to 0 (complete loss of function). Thus, the aggregate score can span from a maximum of 48, indicating full functionality, to a minimum of 0, reflecting total loss of physical function. For a repeatedly measured continuous variable such as ALSFRS-R score, Mixed Models Repeated Measures (MMRM) analyses are typically employed.
@@ 2 False Approval Rate (FAR) in a regulatory setting
For regulatory agencies like the European Medical Agency (EMA) and the United States Food and Drug Administration (FDA), the purpose of Type I error rate control is to control the rate of approving a treatment (e.g., drug) when it is not efficacious, which we refer to as False Approval Rate (FAR) control. To balance patient benefits and risks, the US FDA requires pharmaceutical companies to provide evidence of a new drugs’ safety and efficacy before it is approved for public use, for which the FDA permits a maximum of five incorrect approvals for every 100 ineffective compounds tested, effectively a 5-per-100 rate. To achieve this, the FDA mandates controlling familywise error rate at $5\%$ within each study. In this paper, we refer to each clinical trail as a “study” and we will show that across studies, FAR is controlled, regardless of whether the studies are independent or correlated in families of platform trials.
@@ 2.1 Two layers of families and error rates
Within each study, there may be multiple doses and/or endpoints. The within-study family of inferential statements is the collection of (prespecified) statements to be made about them (e.g. confidence intervals for low dose efficacy in the primary endpoint and high dose efficacy in the secondary endpoint).
@@
In the Healey ALS Platform Trial, within each regimen, there exists a family of efficacy assessment on primary outcomes (e.g., ALSFRS-R total score disease progression) and secondary outcomes (e.g., respiratory function). Subsequently the regimens in the platform trial form a higher layer of family of studies, each of which either succeeds or fails in claiming efficacy of the compound compared to the placebo.
@@ 2.2 Tukey’s three Multiple Comparison error rates
It can perhaps be said that Tukey (1953), which has been reprinted as Tukey (1994), started Multiple Comparisons as a discipline in Statistics. The form of statistical inference that John W. Tukey preferred was confidence interval. He defined the following error rates:
@@
Therefore, instead of the transposed familywise error rate (abbreviated as FWER), we use Tukey’s original ordering error rate familywise (abbreviating it as ERFw). Also contrary to popular perception, it is not true that the only alternative to ERFw is false discovery rate (FDR). While FDR control may be appropriate for medical image reconstruction, it is well known that FDR control is inappropriate for controlling errors within a clinical trial because it can be manipulated by sponsors (see Section 6 of Finner and Roter (2001)). Which multiple comparison error rate to control depends on the specific application as we will discuss below.
2.3 Error Rate Familywise (ERFw) control Guards regulatory False Approval Rate (FAR)
Within a study (a clinical trial), an incorrect statement about a primary endpoint can lead to an incorrect approval of a compound, and an incorrect statement about a secondary endpoint can lead to an incorrect claim of additional efficacy on the drug label. Since any incorrect statement can lead to a (regulatory) incorrect decision, this has led to controlling error rate familywise (ERFw) within each study since the early 2000s.
Different treatment arms within a platform trial may share a common control arm and therefore are not independent. Across the infinitely many studies in independent trials and in platform trials, the following theorem allows us to show if error rate familywise (ERFw) of every study within each platform trial is set at the 5% level, the aforementioned dependence does not affect regulatory false approval rate being no more than 5-per-100 (or 5%), so long as dependence is short-term (within each platform trial only, while studies in different platform trials remain independent).
@@ 2.3 Error Rate Familywise (ERFw) control Guards regulatory False Approval Rate (FAR)
Even though individual studies and those in platform trials are conducted at overlapping times, we can think of them as occurring in a (countably) infinite sequence of clinical trials, indexed as Study $\ldots,-1,0,1,\ldots$. Consider a finite number of $k$ studies, some of which are in platform trials while others are not. Within these $k$ studies, some compounds have efficacy while others do not. Studies with efficacious compounds that fail, i.e. failing to reject the null hypothesis of no-efficacy, do not lead to incorrect regulatory approval. We thus isolate those studies with compounds that lack efficacy (and therefore should not gain regulatory approval) as studies with index in $N$, and let $n$ denote its cardinality (the number of studies with index in $N$).
Let $\ldots,Y_{-1},Y_{0},Y_{1},\ldots$ be the indicator functions of past, present, and future studies that lead to efficacy claims from rejection of the null hypothesis that the compound studied lacks efficacy. Among studies with index in $N$ that should not lead to regulatory approvals, we relabel their $Y_{i}$ as $X_{i}$, i.e., $X_{i}$ is the indicator of a Type I error being committed as a result of Study $i,~{}i\in N,$ a study with compound that lacks efficacy,
@@
Then $E(X_{i})=\mu_{i}$ for all $i\in N$. For $X_{i},X_{j}$ within the same platform trial, the correlation $\rho_{j-i}$ in (2.1) is between zero and one. But with short term dependence, for $X_{i},X_{j}$ in different platform trials, $\rho_{j-i}=0$. For example, if the practical maximum number of study arms within a single Platform trial is 100, then the $\rho_{j-i}$ in (2.1) would satisfy
@@
Let $m$ = $|j-i|$, then $\rho_{m}=0,~{}m=101,102,\ldots$ for platform trials. Therefore the series $\sum_{1}^{\infty}\rho_{i}$ is convergent. The variance convergence condition is also satisfied since $\sigma^{2}_{i}=\mu_{i}(1-\mu_{i})\leq 1$ and $\sum_{i=1}^{\infty}(\log i)^{2}/i^{2}$ is a convergent series.
Let $V$ denote the number of studies incurring Type I errors in $N$, i.e.,
$V=\sum_{i\in N}X_{i}.$
Assuming $P(X_{i}=1)=5\%$ for every $i\in N$, the expectation of $V$ can be expressed as the sum of the probabilities for each $X_{i}$, leading to:
$E[V]$ $=$ $\sum_{i\in N}P(X_{i}=1)=5\%\times n.$ (2.4)
This additive multiplicity adjustment (2.4) for controlling ERpF is an exact equality regardless of the dependency structure among the $X_{i}$ variables, whether they are independent, positively dependent, or negatively dependent. It should in no way be confused with the Bonferroni inequality multiplicity adjustment for the purpose of controlling ERFw.
Moreover, Theorem 2.1 assures that as the number of studies ($n$) grows infinitely large, the proportion of studies incurring Type I errors ($V/n$), i.e., ERpF, will converge to the probability of a Type I error in any single study (i.e., ERFw)—$5\%$ in this case. In other words, by the additive multiplicity adjustment (on line 360 of Food and Drug Administration, (2023)), controlling the Type I error rate familywise (ERFw) of each study exactly at 5% leads to controlling the error rate per family (ERpF) at exactly 5% for a family of $n$ studies with in-efficacious compounds (which we phrase as 5-per-100). More generally, regardless of whether studies in the families are independent or dependent, having an error rate per family no more than $5\%\times n$ for a family of $n$ studies means to expect no more than one incorrect approval per a family of 20 studies, five incorrect approvals per a family of 100 studies, etc. Such precision in error control exemplifies the regulatory commitment to balancing risk and benefit for patients, challenging the perception suggested by Bai et al (2020) Bai et al., (2020).
@@ 3 Stability of FAR and Conditional Assessment
While reassuring that a 5% ERFw within each study results in an ERpF of 5% across all the studies, there is a concern regarding the stability of ERpF across platform trials. Lines 365 and 366 in Section F, “Multiplicity”, of chapter III , “Considerations On Design and Analysis,” in the Food and Drug Administration (2023) guidelines, specifically state that
This section provides a clear demonstration of how to meet this requirement.
Note that it is not whether the unconditional or the conditional error rate should be controlled, as not one error rate control is more meaningful than the other. Both should be assessed because, from elementary statistics, only if one knows the marginal (unconditional) as well as the conditional can one sees the whole picture, i.e. the joint distribution. Thus, think of the platform-specific conditional error rate calculation we describe in this section as within the controlled unconditional error rate over the infinitely many studies over time, conditional on the specific number of studies for a specific platform trial (5 say), with a specific value for the sample mean of the shared control (one SE below its true value say).
@@ 3.1 Conditional Stability of FAR is desirable
An error rate represented as an expectation reflects the average value in the probability distribution of errors made. Although an unconditional error rate might be controlled at $5\%$, actual error rates can vary significantly among different families of studies, depending on the characteristics of studies within each family. For example, some families may experience error rates ranging broadly from $1\%$ to $9\%$, while others might have more stable rates, fluctuating only between $4\%$ and $6\%$. Thus, it’s important to consider not only the unconditional (expected) error rate but also the variability of the actual error rates, conditioned on the specific characteristics of the studies within each family.
@@
Let $N$ represent the subset of index $\{1,\ldots,k\}$ for compounds that lack efficacy. For each $i\in N$, let $Z_{i}$ be the indicator that the $i^{th}$ confidence interval erroneously shows efficacy with $P(Z_{i}=1)=\alpha$. Define $V^{*}$ as the sum of these indicators:
$V^{*}=\sum_{i\in N}Z_{i}.$
The variance of $V^{*}$, given by
$Var(V^{*})=\sum_{i}Var(Z_{i})+\sum_{i\neq j}Cov(Z_{i},Z_{j})$
where $Var(Z_{i})=\alpha(1-\alpha)$, can be calculated once the covariances $Cov(Z_{i},Z_{j})$ for $i\neq j$ are known. This section will discuss two specific concerns on variability of $V^{*}$ associated with sharing a control in a platform trial.
3.1.1 Effect of correlation induced by sharing control on variability of $V^{*}$
In platform trials, there are several methods to compute the covariances of non-coverage events across different treatment arms. Specifically, let $T^{\star}_{i}$ represent the pivoted statistic used to compare the new compound $i$, $i\in N$, against a shared control in a platform trial, such that
has a null distribution free of the efficacy parameter, where $\hat{\mu}_{i}$ and $\hat{\mu}_{SC}$ are estimated mean of the new compound $i$ and the shared control, respectively. For proof of concept, consider the case where each pair $(T^{\star}_{i},T^{\star}_{j})$, with $i\neq j$, follows a bivariate standard Normal distribution. This scenario is particularly applicable in therapeutic areas like Alzheimer’s Disease (AD) and Type 2 Diabetes Mellitus (T2DM), where the primary efficacy endpoints11footnotetext: which excludes the score test version of the logRank test are well-defined and standardized. The covariance between non-coverage indicators $Z_{i}$ and $Z_{j}$ is given by: